Find the first two terms of an arithmetic sequence in which the fourth term is -8 and the 101st term is -396?
Answer by ewatrrr(24785)

a4 = a1+3d = -8
a1 = -8-3d
.......
a101 = (-8-3d) + 100d = -396
97d = 388
d = 4
.........
a1 = -8-3(4) = -20